OpenStudy (anonymous):

2^x+4^(x+5)=4^(2x-1)

OpenStudy (anonymous):

2^x • 4^(x+5) = 4^(2x-1) The trick to solving this is to CHANGE THE BASE to a common one, in this case 2. 2^x • (2²)^(x+5) = (2²)^(2x-1) 2^x • (2)^(2x+10) = (2)^(4x-2) (2)^(3x+10) = (2)^(4x-2) Since the bases are the same, the powers must also be the same. (3x+10) = (4x-2) x = 12

OpenStudy (anonymous):

if this is right can you help me with my problem :D?

OpenStudy (anonymous):

ha well itll take a while to find out, but whats ur problem??

OpenStudy (anonymous):

Which of the following is a false statement? There is only one correct answer for the surface area of a composite figure. There is more than one way to calculate the surface area of a composite figure. The surface area of a composite figure is measured in square units. There is only one way to calculate the surface area of a composite figure.

OpenStudy (anonymous):

=P

OpenStudy (anonymous):

Well now this one is a brain scratcher... hmmm :P lol

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

ha sorry i honestly have no idea lol

OpenStudy (anonymous):

its ok =p no one else did to :(

OpenStudy (anonymous):

I'll ask my dad when he gets home >,< he knows math very well

OpenStudy (anonymous):

What about your question =)?

OpenStudy (anonymous):

wtf? thats math? lol

OpenStudy (anonymous):

yup o-o

OpenStudy (anonymous):

What grade are you in >.<

OpenStudy (anonymous):

I am a senior in hisgh school :P

OpenStudy (anonymous):

=D I just started high school I'm in 7th grade :D

OpenStudy (anonymous):

or middle school o_o idk

OpenStudy (anonymous):

you... dont know? O.o

OpenStudy (anonymous):

I just know I am in 7th grade ._.